# Satellite Autonomous Clock Fault Monitoring with Inter-Satellite Ranges Using Euclidean Distance Matrices

### v2

@@ Appendix A Computing the Test Threshold Using Imhof’s Method
In the rank-decomposition of $\mathbf{F}^{n,m}$, the column matrix has the first column has a one in each fault entry and a zero otherwise. This column spans all the $n-m$ fault-free columns. The remaining $m$ columns correspond to the columns in $\mathbf{F}^{n,m}$ with faults and have a zero at the fault entry. In the example above, the entry $2$ is illustrative. In general, the entry will depend on the fault magnitudes, which may be different. With probability zero, the faults across satellites exactly cancel out and lead to degeneracy in the fault detection, as discussed below. $\mathbf{F}^{n,m}$ achieves full rank when the number of faults reaches $m=n-1$. In general, $\mathrm{rank}(\mathbf{F}^{n,m})=\min(1+m,n)$ if $m>0$ and $\mathrm{rank}(\mathbf{F}^{n,0})=0$. By similar argument, $\mathrm{rank}(\mathbf{F}^{n,m}\circ\mathbf{F}^{n,m})=\mathrm{rank}(\mathbf{F}^{n,m})$ since we will have one column indicating the fault entries and $m$ columns associated with the columns of $\mathbf{F}^{n,m}\circ\mathbf{F}^{n,m}$ with faults, which are the same columns as $\mathbf{F}^{n,m}$. This overall rank relationship often holds even if the faults are of different non-zero magnitudes. However, if the faults cancel out, the rank will drop. For example, in the case above, if $b_{2}=2\bar{f}$ and $b_{3}=b_{4}=-\bar{f}$, then $\mathrm{rank}(\mathbf{F}^{n,m})=3$. Therefore, $\mathrm{rank}(\mathbf{F}^{n,m})\leq\min(1+m,n)$ for arbitrary fault sizes. Nevertheless, the term $\mathrm{rank}(\mathbf{F}^{n,m}\circ\mathbf{F}^{n,m})$ can still be rank $m+1$ even if $\mathrm{rank}(\mathbf{F}^{n,m})<m+1$ since the squaring can remove the linear cancellation.
@@
[…]
@@
First, geometric centering removes the two rank 1 matrices, as shown in Equations (89) and (90) (dokmanic2015euclidean).
@@
For the $-2\mathbf{X}^{\top}\mathbf{X}$ term, notice that the mean of the points is $\mu=\frac{1}{n}\mathbf{X}\mathbf{1}\in\mathbb{R}^{d}$. Using this property yields Equation (91), which is a Gram matrix of the form $\mathbf{X}_{c}^{\top}\mathbf{X}_{c}$ for the point matrix $\mathbf{X}_{c}$ centered about the origin (dokmanic2015euclidean). This shift will not change the rank, meaning $\mathrm{rank}(\mathbf{X}^{\top}\mathbf{X})=\mathrm{rank}(\mathbf{X}_{c}^{\top}\mathbf{X}_{c})$.
$-\frac{1}{2}\mathbf{J}^{n}(-2\mathbf{X}^{\top}\mathbf{X})\mathbf{J}^{n}=(\mathbf{X}-\mu{\mathbf{1}}^{\top})^{\top}(\mathbf{X}-\mu{\mathbf{1}}^{\top})=\mathbf{X}_{c}^{\top}\mathbf{X}_{c}$ (91)
@@ Appendix F The Derivation of the Scale of Chi-squared Distribution
where $\mathbf{\Delta G}=\mathbf{J^{5}}(\mathbf{D}\circ\mathbf{E})\mathbf{J^{5}}$. Here, $\mathbf{G}$ has rank 3, and $\mathbf{\tilde{G}}$ has rank 4, as proved in Corollary 4.4.1. Therefore, $\text{rank}(\mathbf{\Delta G})\geq 4-3=1$, by rank subadditivity. Denote the singular value decomposition of G as
@@
[…]
@@
Since
$\mathbb{E}[\lambda_{4}]$ $=\mathbb{E}\|\mathbf{U_{0}^{\top}\Delta GV_{0}}\|_{F}=\mathbb{E}[\|\mathbf{U_{0}^{\top}J^{5}(D\circ E)J^{5}V_{0}}\|_{F}]=0\quad(\because\mathbb{E}[E_{ij}]=0),$ (99)
the squared 4th singular value $\lambda_{4}^{2}$ follows a scaled chi-squared distribution with degree of freedom = 1. The scale of the distribution is equivalent to its expectation of the mean value, which is
$s$ $=\mathbb{E}\left[\|\mathbf{U_{0}^{\top}J^{5}(D\circ E)J^{5}V_{0}}\|_{F}^{2}\right]$ (100)
@@
Therefore, the scale of the distribution is
$s$ $=\sum_{a=1}^{2}\sum_{b=1}^{2}\left[\sum_{i=1}^{5}\sum_{j=1}^{5}2\sigma_{m}^{2}\left(D_{ij}\hat{U}_{i,a}\hat{V}_{j,b}\right)^{2}\right]$ (102)
$=2\sum_{i=1}^{5}\sum_{j=1}^{5}\left(\sigma_{m}\mathbf{D}_{ij}\right)^{2}\cdot\|\mathbf{\hat{U}}_{i,:}\|^{2}\cdot\|\mathbf{\hat{V}}_{j,:}\|^{2}\quad(\mathbf{\hat{U}=J^{5}U_{0},\hat{V}=J^{5}V_{0}}\in\mathbb{R}^{5\times 2})$
@@
$s$ $=2\sum_{i=1}^{5}\sum_{j=1}^{5}\left(\sigma_{m}\mathbf{\tilde{D}}_{ij}\right)^{2}\cdot\|\mathbf{\hat{U}}_{i,:}\|^{2}\cdot\|\mathbf{\hat{V}}_{j,:}\|^{2}\quad(\mathbf{\hat{U}=J^{5}\tilde{U}_{0},\hat{V}=J^{5}}\tilde{V}_{0}\in\mathbb{R}^{5\times 2})$ (104)
@@ Appendix G The Derivation of the Non-centrality Parameter When Fault Exists
The expectation of the singular value is
$\mathbb{E}[\lambda_{4}]$ $=\mathbb{E}[\|\mathbf{U_{0}}^{T}\mathbf{\Delta GV_{0}}+\mathbf{U_{0}}^{T}\mathbf{\Delta G^{f}V_{0}}\|_{F}]$ (110)
$=\mathbb{E}[\|\mathbf{U_{0}}^{\top}\mathbf{J^{5}}(\mathbf{D}\circ\mathbf{E})\mathbf{J^{5}V_{0}}+\mathbf{U_{0}}^{\top}\mathbf{J^{5}}(\mathbf{D}\circ\mathbf{F})\mathbf{J^{5}V_{0}}\|_{F}]$
$>0\quad\left(\because\mathbb{E}[E_{ij}]=0,\mathbb{E}[(D\circ F)_{ij}]=\begin{cases}D_{ij}f_{ij}\neq 0&i=s_{f}\ \text{or}\ j=s_{f}\\ 0&\text{otherwise}\end{cases}\right)$
Therefore, $\lambda_{4}^{2}$ follows the non-central chi-squared distribution. The expectation of the mean of the squared 4th singular value is
$\mathbb{E}[\lambda_{4}^{2}]$ $=\mathbb{E}\left[\|\mathbf{U_{0}}^{T}\mathbf{\Delta GV_{0}}\|_{F}^{2}\right]+\mathbb{E}\left[\sum_{i=1}^{2}\sum_{j=1}^{2}(\mathbf{U_{0}}^{T}\mathbf{\Delta GV_{0}})_{ij}(\mathbf{U_{0}}^{T}\mathbf{\Delta G^{f}V_{0}})_{ij}\right]+\|\mathbf{U_{0}}^{T}\mathbf{\Delta G^{f}V_{0}}\|_{F}^{2}$ (111)
@@
$=s\left(1+\frac{1}{s}\|\mathbf{U_{0}}^{\top}\mathbf{J}(\mathbf{D}\circ\mathbf{F})\mathbf{JV_{0}}\|_{F}^{2}\right)$
@@
$\lambda_{s_{f}}=\frac{1}{s}\|\mathbf{U_{0}}^{\top}\mathbf{J^{5}}(\mathbf{D}\circ\mathbf{F})\mathbf{J^{5}V_{0}}\|_{F}^{2}$ (112)

### v3

@@ Appendix A Computing the Test Threshold Using Imhof’s Method
In the rank-decomposition of $\mathbf{F}^{n,m}$, the column matrix has the first column has a one in each fault entry and a zero otherwise. This column spans all the nmn-m fault-free columns. The remaining $m$ columns correspond to the columns in $\mathbf{F}^{n,m}$ with faults and have a zero at the fault entry. In the example above, the entry $2$ is illustrative. In general, the entry will depend on the fault magnitudes, which may be different. With probability zero, the faults across satellites exactly cancel out and lead to degeneracy in the fault detection, as discussed below. $\mathbf{F}^{n,m}$ achieves full rank when the number of faults reaches m=n1m=n-1. In general, $\mathrm{rank}(\mathbf{F}^{n,m})=\min(1+m,n)$ if $m>0$ and $\mathrm{rank}(\mathbf{F}^{n,0})=0$. By similar argument, $\mathrm{rank}(\mathbf{F}^{n,m}\circ\mathbf{F}^{n,m})=\mathrm{rank}(\mathbf{F}^{n,m})$ since we will have one column indicating the fault entries and $m$ columns associated with the columns of $\mathbf{F}^{n,m}\circ\mathbf{F}^{n,m}$ with faults, which are the same columns as $\mathbf{F}^{n,m}$. This overall rank relationship often holds even if the faults are of different non-zero magnitudes. However, if the faults cancel out, the rank will drop. For example, in the case above, if $b_{2}=2\bar{f}$ and $b_{3}=b_{4}=-\bar{f}$, then $\mathrm{rank}(\mathbf{F}^{n,m})=3$. Therefore, $\mathrm{rank}(\mathbf{F}^{n,m})\leq\min(1+m,n)$ for arbitrary fault sizes. Nevertheless, the term $\mathrm{rank}(\mathbf{F}^{n,m}\circ\mathbf{F}^{n,m})$ can still be rank $m+1$ even if $\mathrm{rank}(\mathbf{F}^{n,m})<m+1$ since the squaring can remove the linear cancellation.
@@
where $\Delta_{ij}$ are leftover terms from the row-reduction. In the rank-decomposition of $\mathbf{F}^{n,m}\circ\mathbf{S}^{n,d}$, $m$ columns correspond to the columns in $\mathbf{F}^{n,m}$ with faults and have a zero at the fault entry, just as before. However, now, one column is generally not enough to span the fault-free columns since the entries in $\mathbf{S}^{n,d}$ are not linearly related. For example, $(s_{12},s_{13},s_{14})$ is generally not a linear scaling of $(s_{25},s_{35},s_{45})$. These fault-free columns constitute an $m$ dimensional subspace, from which will need $m$ columns to fully span. $\mathbf{F}^{n,m}\circ\mathbf{S}^{n,d}$ achieves full rank when the number of faults reaches $m=\lceil n/2\rceil$, where $\lceil\cdot\rceil$ is the ceiling function. At that point, the subspace of fault-free columns is large enough that the fault columns will contribute to the span. Therefore, $\mathrm{rank}(\mathbf{F}^{n,m}\circ\mathbf{S}^{n,d})\leq\min(2m,n)$ where we do not have equality when $\mathbf{S}^{n,d}$ is degenerate (i.e., if the fault satellites are above or below a plane of fault-free satellites).
@@
First, geometric centering removes the two rank 1 matrices, as shown in Equations (90) and (91) [7].
@@
For the $2\mathbf{X}^{\top}\mathbf{X}$ term, notice that the mean of the points is $\mu=\frac{1}{n}\mathbf{X}\mathbf{1}\in\mathbb{R}^{d}$. Using this property yields Equation (92), which is a Gram matrix of the form $\mathbf{X}_{c}^{\top}\mathbf{X}_{c}$ for the point matrix $\mathbf{X}_{c}$ centered about the origin [7]. This shift will not change the rank, meaning $\mathrm{rank}(\mathbf{X}^{\top}\mathbf{X})=\mathrm{rank}(\mathbf{X}_{c}^{\top}\mathbf{X}_{c})$.
$\frac{1}{2}\mathbf{J}^{n}(-2\mathbf{X}^{\top}\mathbf{X})\mathbf{J}^{n}=(\mathbf{X}-\mu{\mathbf{1}}^{\top})^{\top}(\mathbf{X}-\mu{\mathbf{1}}^{\top})=\mathbf{X}_{c}^{\top}\mathbf{X}_{c}$ (92)
@@ Appendix F The Derivation of the Scale of Chi-squared Distribution
where $\mathbf{\Delta G}=\mathbf{J^{5}}(\mathbf{D}\circ\mathbf{E})\mathbf{J^{5}}$. The second-order error term $(\mathbf{E}\circ\mathbf{E})$ is approximated as zero in the third line. Here, $\mathbf{G}$ has rank 3, and $\mathbf{\tilde{G}}$ has rank 4, as proved in Corollary 4.4.1. Therefore, $\text{rank}(\mathbf{\Delta G})\geq 4-3=1$, by rank subadditivity. Denote the singular value decomposition of G as
@@
$\approx\mathbf{\Sigma+U^{\top}(\Delta G)V+\Sigma V^{\top}(\Delta V)+\Delta U^{\top}U\Sigma}\quad(\because\text{2nd and 3rd order perturbation }\approx 0)$
@@
We observe that the expectation of the matrix elements inside the norm is zero:
$\mathbb{E}[(\mathbf{U}_{0}^{\top}\Delta\mathbf{G}\mathbf{V}_{0})_{ij}]=\left(\mathbf{U}_{0}^{\top}\mathbf{J}^{5}(\mathbf{D}\circ\mathbb{E}[\mathbf{E}])\mathbf{J}^{5}\mathbf{V}_{0}\right)_{ij}=0\quad(\because\mathbb{E}[E_{ij}]=0).$ (100)
Moreover, the term $(\mathbf{U}_{0}^{\top}\Delta\mathbf{G}\mathbf{V}_{0})_{ij}$ is a linear combination of the elements of $\mathbf{E}$, where $E_{ij}$ follows a Gaussian distribution. Since a linear combination of Gaussian variables is also Gaussian, the term itself is normally distributed. Combined with the zero-mean property, the squared 4th singular value $\lambda_{4}^{2}$ follows a scaled central chi-squared distribution with 1 degree of freedom ($k$), where $\lambda_{4}/s^{2}$ follows a chi-squared distribution. Since the mean of a chi-squared distribution can be written as $ks^{2}$ (k=1), $s^{2}$ is given by
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{2}}$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{E}[\lambda_{4}^{2}]}$ (101)
$=\mathbb{E}\left[\|\mathbf{U_{0}^{\top}J^{5}(D\circ E)J^{5}V_{0}}\|_{F}^{2}\right]$
@@
Therefore, the squared scale of $\lambda_{4}$ is
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{2}}$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum_{a=1}^{2}\sum_{b=1}^{2}\sum_{i=1}^{5}\sum_{j=1}^{5}\sigma_{m}^{2}D_{ij}^{2}\left(\hat{U}_{i,a}\hat{V}_{j,b}+\hat{U}_{j,a}\hat{V}_{i,b}\right)^{2}}$ (103)
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{2}}$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum_{a=1}^{2}\sum_{b=1}^{2}\sum_{i=1}^{5}\sum_{j=1}^{5}\left(\sigma_{m}\mathbf{\tilde{D}}_{ij}\right)^{2}\cdot\left(\hat{U}_{i,a}\hat{V}_{j,b}+\hat{U}_{j,a}\hat{V}_{i,b}\right)^{2}}\quad(\mathbf{\hat{U}=J^{5}\tilde{U}_{0},\hat{V}=J^{5}\tilde{V}_{0}}\in\mathbb{R}^{5\times 2})$ (105)
@@ Appendix G The Derivation of the Non-centrality Parameter When Fault Exists
We analyze the expectation of the matrix term inside the norm:
$\mathbb{E}[\mathbf{U}_{0}^{\top}(\mathbf{\Delta G}+\mathbf{\Delta G}^{f})\mathbf{V}_{0}]$ $=\mathbb{E}[\mathbf{U}_{0}^{\top}\mathbf{\Delta G}\mathbf{V}_{0}]+\mathbf{U}_{0}^{\top}\mathbf{\Delta G}^{f}\mathbf{V}_{0}$ (111)
$=\mathbf{0}+\mathbf{U}_{0}^{\top}\mathbf{J}^{5}(\mathbf{D}\circ\mathbf{F})\mathbf{J}^{5}\mathbf{V}_{0}$
$\neq\mathbf{0}\quad(\because\mathbf{D}\circ\mathbf{F}\neq\mathbf{0})$
The underlying variable is the sum of a zero-mean Gaussian term ($\mathbf{U}_{0}^{\top}\mathbf{\Delta G}\mathbf{V}_{0}$) and a non-zero constant bias term ($\mathbf{U}_{0}^{\top}\mathbf{\Delta G}^{f}\mathbf{V}_{0}$). Because the underlying variable is a Gaussian with a non-zero mean, its squared Frobenius norm $\lambda_{4}^{2}$ follows a non-central chi-squared distribution.
The mean of this non-central chi-squared distribution is
$s^{2}(k+\lambda)=\mathbb{E}[\lambda_{4}^{2}]$ $=\mathbb{E}\left[\|\mathbf{U_{0}}^{T}\mathbf{\Delta GV_{0}}\|_{F}^{2}\right]+\mathbb{E}\left[\sum_{i=1}^{2}\sum_{j=1}^{2}(\mathbf{U_{0}}^{T}\mathbf{\Delta GV_{0}})_{ij}(\mathbf{U_{0}}^{T}\mathbf{\Delta G^{f}V_{0}})_{ij}\right]+\|\mathbf{U_{0}}^{T}\mathbf{\Delta G^{f}V_{0}}\|_{F}^{2}$ (112)
@@
$={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{2}}\left(1+\frac{1}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{2}}}\|\mathbf{U_{0}}^{\top}\mathbf{J}(\mathbf{D}\circ\mathbf{F})\mathbf{JV_{0}}\|_{F}^{2}\right)$
where $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{2}$} is the squared scale given in (105)
@@
$\lambda_{s_{f}}=\frac{1}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s^{2}}}\|\mathbf{U_{0}}^{\top}\mathbf{J^{5}}(\mathbf{D}\circ\mathbf{F})\mathbf{J^{5}V_{0}}\|_{F}^{2}$ (113)
Appendix H Generalization of the Test Statistic Distribution for $n\geq 5$
We generalize the derivation from Appendix F to an arbitrary number of satellites $n\geq 5$. We define the generalized test statistic $\gamma_{test}$ as the total energy of the noise subspace singular values:
$\gamma_{test}=\sum_{k=4}^{n-1}\tilde{\lambda}_{k}^{2}$ (114)
where $\tilde{\lambda}_{k}$ are the singular values of the observed (noisy) GCEDM $\tilde{G}$. This statistic is equivalent to the squared Frobenius norm of the matrix projected onto the active noise subspace.
Since the true noiseless matrix $G$ is unknown, we approximate the active noise subspace using the singular vectors of the measured matrix $\tilde{G}$. Let the SVD of the observed matrix be $\tilde{G}=\tilde{U}\tilde{\Sigma}\tilde{U}^{\top}$. We partition the singular vectors $\tilde{U}$ as follows:
$\tilde{U}=\begin{bmatrix}\underbrace{\tilde{u}_{1}\quad\tilde{u}_{2}\quad\tilde{u}_{3}}_{\tilde{U}_{S}}&\underbrace{\tilde{u}_{4}\quad\ldots\quad\tilde{u}_{n-1}}_{\tilde{U}_{noise}}&\underbrace{\tilde{u}_{n}}_{\tilde{u}_{\mathbf{1}}}\end{bmatrix}$ (115)
Here, $\tilde{U}_{S}$ spans the signal subspace. The vectors $\tilde{u}_{4},\ldots,\tilde{u}_{n-1}$ span the active noise subspace. The final vector $\tilde{u}_{n}$ corresponds to the structural null direction parallel to $\mathbf{1}$. Due to geometric centering ($\tilde{G}\mathbf{1}=\mathbf{0}$), the singular value associated with $\tilde{u}_{n}$ remains strictly zero.
We approximate the projection of the noise perturbation $\Delta G$ using this observed basis $\tilde{U}_{noise}$. Let $M_{noise}\approx\tilde{U}_{noise}^{\top}\Delta G\tilde{U}_{noise}$. Since $\Delta G$ is linear in the Gaussian measurement noise $E$, $M_{noise}$ is a symmetric matrix of zero-mean Gaussian variables. The sum of squares of its independent elements follows a Chi-squared distribution with degrees of freedom $k$:
$k=\frac{(n-4)(n-3)}{2}$ (116)
(e.g., $k=1$ for $n=5$). The scaled test statistic follows $\frac{\gamma_{test}}{s^{2}}\sim\chi^{2}(k,0)$, where the squared scale parameter $s^{2}$ is computed using the observed singular vectors $\tilde{u}_{a}$ (columns of $\tilde{U}_{noise}$) and the observed distances $\tilde{D}_{ij}$:
$s^{2}\approx\frac{1}{k}\sum_{a=4}^{n-1}\sum_{b=4}^{n-1}\left(\sum_{i=1}^{n}\sum_{j=1}^{n}(\sigma_{m}\tilde{D}_{ij})^{2}(\tilde{u}_{i,a}\tilde{u}_{j,b}+\tilde{u}_{j,a}\tilde{u}_{i,b})^{2}\right)$ (117)
This approximation holds because the perturbation of the singular vectors corresponding to the noise-subspace is small (Equation (98)), allowing $\tilde{U}_{noise}$ (obtained from $\mathbf{\tilde{G}}$) to serve as a valid approximation for the true noise subspace $\hat{U}_{N}$ (of $\mathbf{G}$) in the variance computation. Note that for $n=5$, equation (117) reduces to the result derived in Equation (105).
When a fault occurs, the GCEDM includes a deterministic bias term $\Delta G^{f}=J^{n}(D\circ F)J^{n}$. The projection onto the noise subspace becomes $M_{total}=M_{noise}+M_{bias}$, where $M_{bias}=\hat{U}_{N}^{\top}\Delta G^{f}\hat{U}_{N}$. The test statistic $\gamma_{test}\approx\|M_{total}\|_{F}^{2}$ then follows a non-central Chi-squared distribution:
$\frac{\gamma_{test}}{s^{2}}\sim\chi^{2}(k,\lambda_{nc})$ (118)
The non-centrality parameter $\lambda_{nc}$ can be computed as:
$\lambda_{nc}=\frac{1}{s^{2}}\|M_{bias}\|_{F}^{2}=\frac{1}{s^{2}}\sum_{a=1}^{n-4}\sum_{b=1}^{n-4}(\hat{u}_{a}^{\top}\Delta G^{f}\hat{u}_{b})^{2}$ (119)
For $n=5$, this reduces to the result $\lambda_{s_{f}}=\frac{1}{s^{2}}(u_{0}^{\top}\Delta G^{f}u_{0})^{2}$ derived in Equation (113).
Appendix I Proof of Proposition 4.5
Proof.
We assume the graph $G$ is redundantly rigid. A graph is fault disprovable if the injection of a fault bias makes the set of edge constraints inconsistent (i.e., the graph cannot be embedded in Euclidean space).
Let $v_{f}\in V$ be a satellite experiencing a clock fault. This fault introduces a bias $b\neq 0$ to the measured ranges. The weights of the edges incident to $v_{f}$ are given by:
$W(u,v_{f})=\begin{cases}\|\mathbf{x}_{u}-\mathbf{x}_{v_{f}}\|+b&\text{if }(u,v_{f})\in E_{m}\\ \|\mathbf{x}_{u}-\mathbf{x}_{v_{f}}\|&\text{if }(u,v_{f})\in E_{e}\end{cases}$ (120)
where the measurement noise and ephemeris error are not considered in this proof.
Case 1: Sufficient Condition. Assume $v_{f}$ has at least one incident edge $e_{meas}=(u,v_{f})\in E_{m}$. Since $G$ is redundantly rigid, the distance between $u$ and $v_{f}$ is structurally constrained by the remaining edges in $E\setminus\{e_{meas}\}$. The faulty weight $W(e_{meas})$ includes the bias $b$, creating a geometric contradiction with the constraints imposed by the rest of the graph. Thus, the graph becomes unrealizable, and the fault is detectable.
Case 2: Necessary Condition. Assume $v_{f}$ has no incident edges in $E_{m}$ (i.e., $\text{deg}_{E_{m}}(v_{f})=0$). In this case, all edges connected to $v_{f}$ are from the set $E_{e}$. The weights of these edges are derived solely from the ephemeris, which is independent of the instantaneous clock jump fault. Consequently, the geometry of $v_{f}$ in the graph remains consistent with the ephemeris model, regardless of the actual clock fault. The graph remains realizable, and the fault cannot be detected by analyzing the rigidity or embeddability of $G$.
Therefore, for the graph to be fault-disprovable for any single satellite fault, it is necessary that every node $v\in V$ has at least one incident edge in $E_{m}$. ∎